OpenStudy (anonymous):

***medals*** Work-Out Corner has 5 more than 3 times as many exercise bicycles as The Gym. Together they have 21 bicycles. Solve the equation x + 3x + 5 = 21 to find the number of bicycles at Work-Out Corner. A. 4 bicycles B. 17 bicycles C. 7 bicycles D. 25 bicycles

OpenStudy (anonymous):

@Fanduekisses

OpenStudy (fanduekisses):

A

OpenStudy (fanduekisses):

combine like terms first (x+3x)= 4x +5=21 subtract 5 to both sides= 2x=16, divide 3 to both sides...

OpenStudy (anonymous):

ok thanks, i got one more

OpenStudy (fanduekisses):

B

OpenStudy (fanduekisses):

7.25+x=443.75, subtract 7.25 to both sides, you get x=436.5, but they tell you 25 students went on the trip so you divide that by 25 and get 17.46 ;)

OpenStudy (anonymous):

thank you ;p

OpenStudy (fanduekisses):

You are welcome ;)

OpenStudy (anonymous):

^.^